OpenStudy (tatumlee):

Exponents and Exponential Functions... does anyone have any tips/resources/notes that could help with this topic? Thank you (:

OpenStudy (tatumlee):

@TheSmartOne @pooja195

OpenStudy (sleepyjess):

What specifically do you need help with? All of exponential function? Graphing?

OpenStudy (sleepyjess):

Here's a link that may be helpful, http://www.regentsprep.org/regents/math/algtrig/ATP8b/exponentialfunction.htm

OpenStudy (tatumlee):

Multiplying Powers with the Same Base, Division Properties of Exponents, Exponential Functions... these are the things I struggled with the most, I was wondering if there was any notes or things that might help

OpenStudy (sleepyjess):

Could you give an example of multiplying powers with the same base? I'm not quite sure what that is.

OpenStudy (sleepyjess):

Like \(5^3*5^5\)?

OpenStudy (tatumlee):

Yes I think that is it

OpenStudy (sleepyjess):

Okay, so for multiplying exponents with the same base, the base stays the same and the exponents are added together. For \(5^3*5^5\), that would be \(5^8\)

OpenStudy (bibby):

properties of exponents \(a^0=1\\ a^{-m}=\dfrac{1}{a^m}\\ \dfrac{a^M}{a^N}=a^{(M-N)}\\ a^M*a^N=a^{(M+N)}\\ {(a^M)}^N=a^{M*N}\\ (\dfrac{a}{b})^M=\dfrac{a^M}{b^M}\\ a^M \times b^M = (ab)^M\\ \sqrt[m]{n}=n^{\dfrac{1}{m}}\)

OpenStudy (tatumlee):

I have been trying to find notes like this, I will write them down. I didn't get the notes for this chapter that's why I was asking, thank you guys so much <3

OpenStudy (tatumlee):

And the multiplying powers of the same base included scientific notation, it was confusing

OpenStudy (bibby):

you can post some specific problems if you want it walked through

OpenStudy (tatumlee):

hopefully the notes will be enough for this chapter, if not I will sure for post specific questions, thank you again (: